# Conformal prediction for causal effects of continuous treatments

### v2

@@ Abstract
Uncertainty quantification of causal effects is crucial for safety-critical applications such as personalized medicine. A powerful approach for this is conformal prediction, which has several practical benefits due to model-agnostic finite-sample guarantees. Yet, existing methods for conformal prediction of causal effects are limited to binary/discrete treatments and make highly restrictive assumptions such as known propensity scores. In this work, we provide a novel conformal prediction method for potential outcomes of continuous treatments. We account for the additional uncertainty introduced through propensity estimation so that our conformal prediction intervals are valid even if the propensity score is unknown. Our contributions are three-fold: (1) We derive finite-sample prediction intervals for potential outcomes of continuous treatments. (2) We provide an algorithm for calculating the derived intervals. (3) We demonstrate the effectiveness of the conformal prediction intervals in experiments on synthetic and real-world datasets. To the best of our knowledge, we are the first to propose conformal prediction for continuous treatments when the propensity score is unknown and must be estimated from data.
@@ 1 Introduction
Machine learning (ML) for estimating causal quantities such as causal effects and the potential outcomes of treatments is nowadays widely used in real-world applications such as personalized medicine (Feuerriegel et al., 2024). However, existing methods from causal ML typically focus on point estimates (e.g., Nie et al., 2021; Patrick Schwab et al., 2020), which means that the uncertainty in the predictions is neglected and hinders the use of causal ML in safety-critical applications (Feuerriegel et al., 2024; Kneib et al., 2023). As the following example shows, uncertainty quantification (UQ) of causal quantities is crucial for reliable decision-making.
@@
A powerful method for UQ is conformal prediction (CP) (Lei & Wasserman, 2014; Papadopoulos, 2002; Vovk et al., 2005). CP provides model-agnostic and distribution-free, finite-sample predictive intervals for quantifying uncertainty. CP has been widely used for traditional, predictive ML (e.g., Angelopoulos et al., 2024; Barber et al., 2023; Gibbs et al., 2023), where it has been shown to yield reliable prediction intervals in finite-sample settings (see Fig. 1). Recently, there have been works that adapt CP for estimating causal quantities (see Fig. 2 for an overview). Yet, existing methods for CP focus on binary or discrete treatments (e.g., Alaa et al., 2023; Jonkers et al., 2024; Lei & Candès, 2021), but not continuous treatments, which is our novelty.
@@
However, adapting CP to causal quantities with continuous treatments is non-trivial for two main reasons. Challengea: Intervening on the treatment induces a shift in the covariate distribution, specifically, in the propensity score. As a result, the so-called exchangeability assumption, which is inherent to CP (Vovk et al., 2005), is violated between the observational and interventional distribution, and because of this, standard CP intervals are not valid. Thus, we must later account for the distribution shift and derive treatment-conditional guarantees. Challengeb: Assessing the aforementioned shift in the distribution requires information about the propensity score; yet, the propensity score is typically unknown. Hence, estimating the propensity score introduces additional uncertainty. However, incorporating the additional uncertainty in the overall CP intervals cannot be done in a simple plug-in manner, and it is highly non-trivial.
In this paper, we develop a novel CP method for causal quantities, such as causal effects and potential outcomes of continuous treatments. Our method is carefully designed to account for the additional uncertainty introduced during propensity estimation and is therefore applicable to settings where the propensity score is known or even unknown.
@@ 2 Related Work
UQ for causal effects: Existing methods for UQ of causal quantities are often based on Bayesian methods (e.g., Alaa & van der Schaar, 2017; Hess et al., 2024; Hill, 2011; Jesson et al., 2020). However, Bayesian methods require the specification of a prior distribution based on domain knowledge and are thus neither robust to model misspecification nor generalizable to model-agnostic machine learning models. A common ad hoc method for computing uncertainty intervals is Monte Carlo (MC) dropout (Gal & Ghahramani, 2016). However, MC dropout yields approximations of the posterior distribution, which are not faithful (Le Folgoc et al., 2021).
@@
Conformal prediction: CP (Lei & Wasserman, 2014; Papadopoulos, 2002; Vovk et al., 2005) has recently received large attention for finite-sample UQ. For a prediction model $\phi$ trained on dataset $D_{T}=(X_{i},Y_{i})_{i=1,\ldots,m}$ and a new test sample $X_{k}$, CP aims to construct a prediction interval $C(X_{k})$ such that $P(Y_{k}\in C(X_{k}))\geq 1-\alpha$ for some significance level $\alpha$. We refer to (Angelopoulos et al., 2024) for an in-depth overview. Due to its strong finite-sample mathematical guarantees, CP is widely used for traditional, predictive ML with widespread applications such as in medical settings (Zhan et al., 2020) or drug discovery (Alvarsson et al., 2021; Eklund et al., 2015).
@@
Several extensions have been developed for CP. One literature stream focuses on CP with marginal coverage under distribution shifts between training and test data (e.g., Cauchois et al., 2020; Fannjiang et al., 2022; Gendler et al., 2022; Ghosh et al., 2023; Gibbs & Candès, 2021; Gibbs et al., 2023; Guan, 2023; Lei & Candès, 2021; Podkopaev & Ramdas, 2021; Tibshirani et al., 2019; Yang et al., 2024). Our setting later also involves a distribution shift due to the intervention on the treatment but differs from the latter in that the true distribution shift is unknown. Another literature stream constructs intervals conditional on the variables following the shifted distribution. Since, in general, exact conditional coverage has been proven impossible (Lei & Wasserman, 2014; Vovk, 2012), the works in this literature stream have two key limitations: (1) they only guarantee approximate conditional coverage (e.g., Barber et al., 2021; Cai et al., 2014; Lei & Wasserman, 2014; Romano et al., 2020); or (2) they are restricted to specific data structures such as binary variables (e.g., Lei & Wasserman, 2014; Vovk, 2012). Because of that, none of the existing methods for marginal and conditional coverage can be applied to derive exact finite-sample prediction intervals for causal quantities of continuous treatments.
@@
Conformal prediction for causal quantities: Only a few works focus on CP for causal quantities (see Fig. 2). Examples are methods aimed at off-policy learning (Taufiq et al., 2022; Zhang et al., 2023), conformal sensitivity analysis (Yin et al., 2022), or meta-learners for the conditional average treatment effect (CATE) (Alaa et al., 2023; Jonkers et al., 2024; Wang et al., 2024). However, there are crucial differences to our setting: First, the existing works (a) assume that the propensity is known and thus achieve exact finite-sample coverage guarantees, or the existing works (b) focus on the easier task of giving asymptotic guarantees but then might suffer from under-coverage because of which the intervals are not faithful. On top of that, the existing CP methods are designed for binary or discrete treatments. Applying such methods to discretized continuous treatments leads to ill-defined causal estimands. Therefore, none of the existing methods are applicable to our continuous treatment setting. We offer a detailed discussion in Supplement D.
Figure 2: Key works on CP in causal inference.
@@
Research gap: To the best of our knowledge, no work has provided exact finite-sample prediction intervals for causal quantities of continuous treatments, which is our novelty.
@@ 3 Problem formulation
Throughout this work, we build upon the potential outcomes framework (Rubin, 2005). We denote the potential outcomes after a hard intervention $a^{*}$ by $Y(a^{*})$ and after a soft intervention $A^{*}(x)\sim\tilde{\pi}(a\mid x)=P_{A^{*}\mid X=x}$ by $Y(A^{*}(x))$. [^2] Interventions are characterized by two classes: hard (structural) and soft (parametric) interventions. Hard interventions directly affect the treatment by setting it to a specific value and removing the edge in the graph (as in the do-operator). Soft interventions do not change the structure of the graph but affect the conditional distribution of the treatment given the confounders, i.e., the propensity score. We make three standard identifiability assumptions for causal effect estimation: positivity, consistency, and unconfoundedness (e.g., Alaa et al., 2023; Jonkers et al., 2024). Finally, we consider an arbitrary machine learning model $\phi$ to predict the potential outcomes. Hence, we define the outcome prediction function as $\phi:\mathcal{X}\times\mathcal{A}\rightarrow\mathbb{R}$, $\phi(X,A)\mapsto Y$.
Our objective: In this work, we aim to derive conformal prediction intervals $C(X_{n+1},\Diamond)$ for the prediction of a potential outcome $Y_{n+1}(\Diamond)$ of a new data point under either hard, $\Diamond=a^{*}$, or soft intervention, $\Diamond=A^{*}(X_{n+1})\sim\tilde{\pi}(a\mid X_{n+1})$. The derived intervals are called valid for any new exchangeable sample $X_{n+1}$ with non-exchangeable intervention $\Diamond$, i.e.,
$P(Y_{n+1}(\Diamond)\in C(X_{n+1},\Diamond))\geq 1-\alpha,\quad\Diamond\in\{a^{*},A^{*}(X_{n+1})\}.$ (1)
for some significance level $\alpha\in(0,1)$. Of note, our CP method can later be used with an arbitrary machine learning model $\phi$ for predicting the potential outcomes.
In CP, the interval $C$ is constructed based on so-called non-conformity scores (Vovk et al., 2005), which capture the performance of the prediction model $\phi$. For example, a common choice for the non-conformity score is the residual of the fitted model $s(X,A,Y)=\left|Y-\phi(X,A)\right|$, which we will use throughout our work. For ease of notation, we define $S_{i}:=s(X_{i},A_{i},Y_{i})$.
@@
Second, the propensity score $\pi$ is commonly unknown in observational data and, therefore, must be estimated, which introduces additional uncertainty that one must account for when constructing CP intervals ($\rightarrow$ Challenge  b). Crucially, existing coverage guarantees (e.g., Vovk, 2012; Tibshirani et al., 2019) do not hold in our setting. Instead, we must derive new intervals with valid coverage under propensity shift.
In the following section, we address the above propensity shift by performing a calibration conditional on the propensity shift induced by the intervention, which allows us then to yield valid prediction intervals with significance level $(1-\alpha)$ for potential outcomes of a specific hard or soft intervention. We derive the potential outcomes and emphasize that the extension to causal effects is straightforward. For the latter, one combines the intervals for each potential outcome under a certain treatment and without treatment, so that eventually arrives at CP intervals for the individual treatment effect (ITE). Details are Supplement A.
@@ 4 CP intervals for potential outcomes of continuous treatments
Recall that intervening on test data samples breaks the exchangeability assumption necessary for the validity, i.e., the guaranteed coverage of at least $(1-\alpha)$, of CP intervals. Therefore, we now construct CP intervals where we account for a (potentially unknown) covariate shift in the test data induced by the intervention.
@@
    Example: Imagine a doctor is about to prescribe a medication to a new patient. Instead of prescribing the same dosage as he would have prescribed to a similar patient in the past, the doctor is interested in the potential health outcome of the patient when increasing (or decreasing) the original dosage by amount $\Delta_{A}$.
@@
    Unknown propensity score (see Section  4.2): In observational data, the propensity score is typically unknown. Therefore, we are interested in the effect of hard interventions, i.e., $a^{\ast}$. Here, we face additional uncertainty from the propensity score estimation ($\rightarrow$  Challenge  b).
    Example: In our running example, a patient comes to a new doctor who has never prescribed the respective medication and thus will base the decision on observational data (electronic health records) from other physicians, yet the observational data was collected under a different, unknown treatment policy. Therefore, the prescribed intervention (and thus dosage) cannot be expressed in terms of the policy in the observational data.
@@
Second, our CP method will build upon ideas from so-called split conformal prediction (Papadopoulos, 2002; Vovk et al., 2005), yet with crucial differences. In our methods, the calibration step differs from the standard procedure in that we conditionally calibrate the non-conformity scores depending on the tilting function $f$ to achieve marginal coverage for the interventional – and thus shifted – data.
High-level outline: Our derivation in Sections 4.1 and 4.2 proceed as follows. Following (Gibbs et al., 2023; Romano et al., 2019), we reformulate split conformal prediction as an augmented quantile regression. Let $S_{i}$ represent the non-conformity score of the sample $(X_{i},A_{i},Y_{i})$ for $i=m+1,\ldots,n$ of the calibration dataset and $S_{n+1}=S$ an imputed value for the unknown score of the new sample. We define
$\hat{\theta}_{S}:=\operatorname*{arg\,min}\limits_{\theta\in\mathbb{R}}\frac{1}{n-m}\left(\sum_{i=m+1}^{n}l_{\alpha}(\theta,S_{i})+l_{\alpha}(\theta,S)\right),$ (3)
@@
Of note, $\hat{\theta}_{S}$ is an estimator of the $(1-\alpha)$-quantile of the non-conformity scores (Koenker & Bassett, 1978; Steinwart & Christmann, 2011). Using $\hat{\theta}_{S}$, we then construct the CP interval with the desired coverage guarantee $(1-\alpha)$. However, the interval is only valid for exchangeable data. Quantile regression might yield non-unique solutions, so we later restrict the analysis to solvers invariant to the data ordering.

### v3

@@ Abstract
Uncertainty quantification of causal effects is crucial for safety-critical applications such as personalized medicine. A powerful approach for this is conformal prediction, which has several practical benefits due to model-agnostic finite-sample guarantees. Yet, existing methods for conformal prediction of causal effects are limited to binary/discrete treatments and make highly restrictive assumptions, such as known propensity scores. In this work, we provide a novel conformal prediction method for potential outcomes of continuous treatments. We account for the additional uncertainty introduced through propensity estimation so that our conformal prediction intervals are valid even if the propensity score is unknown. Our contributions are three-fold: (1) We derive finite-sample validity guarantees for prediction intervals of potential outcomes of continuous treatments. (2) We provide an algorithm for calculating the derived intervals. (3) We demonstrate the effectiveness of the conformal prediction intervals in experiments on synthetic and real-world datasets. To the best of our knowledge, we are the first to propose conformal prediction for continuous treatments when the propensity score is unknown and must be estimated from data.
@@ 1 Introduction
Machine learning (ML) for estimating causal quantities such as causal effects and the potential outcomes of treatments is nowadays widely used in real-world applications such as personalized medicine [12]. However, existing methods from causal ML typically focus on point estimates [e.g., 39, 41], which means that the uncertainty in the predictions is neglected and hinders the use of causal ML in safety-critical applications [12, 31]. As the following example shows, uncertainty quantification (UQ) of causal quantities is crucial for reliable decision-making.
@@
A powerful method for UQ is conformal prediction (CP) [35, 40, 53]. CP provides model-agnostic and distribution-free, finite-sample validity guarantees for quantifying uncertainty. CP has been widely used for traditional, predictive ML [e.g., 4, 6, 18], where it has been shown to yield reliable prediction intervals in finite-sample settings (see Fig. 1). Recently, there have been works that adapt CP for estimating causal quantities (see Fig. 2 for an overview). Yet, existing methods for CP focus on binary or discrete treatments [e.g., 2, 27, 36], but not continuous treatments, which is our novelty.
Adapting CP to causal quantities is non-trivial for two main reasons. Challengea: Intervening on the treatment induces a shift in the covariate distribution, specifically in the propensity score. As a result, the so-called exchangeability assumption, which is inherent to CP [53], is violated between the observational and interventional distribution [^1] By interventional we refer to the distribution after the intervention, and because of this, standard CP intervals are not valid. Thus, we must later account for the distribution shift and derive treatment-conditional guarantees. Challengeb: Assessing the aforementioned shift in the distribution requires information about the propensity score; yet, the propensity score is typically unknown. Hence, estimating the propensity score introduces additional uncertainty. However, incorporating the additional uncertainty in the overall CP intervals cannot be done in a simple plug-in manner, and it is highly non-trivial.
Unique to CP for effects of continuous treatments is a third challenge  c: data points with the same treatment value are rarely observed. Thus, we later employ smoothing to model the propensity shift.
In this paper, we develop a CP method for causal quantities, such as potential outcomes, of continuous treatments. Our method is designed to account for the additional uncertainty introduced during propensity estimation and is thus applicable to settings where the propensity score is unknown.
@@ 2 Related Work
UQ for causal effects: Existing methods for UQ of causal quantities are often based on Bayesian methods [e.g., 1, 21, 22, 24]. However, Bayesian methods require the specification of a prior distribution based on domain knowledge and are thus neither robust to model misspecification nor generalizable to model-agnostic machine learning models. A common ad hoc method for computing uncertainty intervals is Monte Carlo (MC) dropout [14]. However, MC dropout yields approximations of the posterior distribution, which are not faithful [34].
@@
Conformal prediction: CP [35, 40, 53] has recently received large attention for finite-sample UQ. For a prediction model $\phi$ trained on dataset $D_{T}=(X_{i},Y_{i})_{i=1,\ldots,m}$ and a new test sample $X_{k}$, CP aims to construct a prediction interval $C(X_{k})$ such that $P(Y_{k}\in C(X_{k}))\geq 1-\alpha$ for some significance level $\alpha$. We refer to [4] for an in-depth overview. Due to its strong finite-sample validity guarantees, CP is widely used for traditional, predictive ML with widespread applications such as in medical settings [59] or drug discovery [3, 10].
@@
Several extensions have been developed for CP. One literature stream focuses on CP with marginal coverage under distribution shifts between training and test data [e.g., 8, 11, 15, 16, 17, 18, 19, 36, 42, 50, 57]. Our setting later also involves a distribution shift due to the intervention on the treatment, but differs from the latter in that the true distribution shift is unknown. Another literature stream constructs intervals conditional on the variables following the shifted distribution. Since, in general, exact conditional coverage has been proven impossible [35, 52], the works in this literature stream have two key limitations: (1) they only guarantee approximate conditional coverage [e.g., 5, 7, 35, 44]; or (2) they are restricted to specific data structures such as binary variables [e.g., 35, 52]. Because of that, none of the existing methods for marginal and conditional coverage can be applied to derive prediction intervals with finite-sample validity guarantees for causal quantities of continuous treatments.
Figure 2: Key works on causal CP.
@@
Conformal prediction for causal quantities: Only a few works focus on CP for causal quantities (see Fig. 2). Examples are methods aimed at off-policy learning [49, 60], conformal sensitivity analysis [58], or meta-learners for the conditional average treatment effect (CATE) [2, 27, 36, 54]. However, there are crucial differences to our setting: First, the existing works (a) assume that the propensity is known and thus achieve finite-sample coverage guarantees, or the existing works (b) focus on the easier task of giving asymptotic guarantees but then might suffer from under-coverage because of which the intervals are not faithful. Only Lei and Candès [36] provides finite-sample coverage guarantees under estimated propensity scores. However, all existing CP methods are designed for binary or discrete treatments. Applying such methods to discretized continuous treatments leads to ill-defined causal estimands. Therefore, none of the existing methods are applicable to our continuous treatment setting. We offer a detailed discussion in Supplement E.
@@
Research gap: To the best of our knowledge, no work has provided prediction intervals with finite-sample validity guarantees for causal quantities of continuous treatments.
@@ 3 Problem formulation
Throughout this work, we build upon the potential outcomes framework [45]. We denote the potential outcomes of a hard intervention $a^{*}$ by $Y(a^{*})$ and of a soft intervention $A^{*}(x)\sim\tilde{\pi}(a\mid x)=P_{A^{*}\mid X=x}$ by $Y(A^{*}(x))$. [^3] Interventions are characterized by two classes: hard (structural) and soft (parametric) interventions. Hard interventions directly affect the treatment by setting it to a specific value and removing the edge in the graph (as in the do-operator). Soft interventions do not change the structure of the graph but affect the conditional distribution of the treatment given the confounders. All interventions affect the propensity score, but the mathematical consequences are different. For soft interventions, the new (interventional) propensity is a function of the original propensity. For hard interventions, the new propensity is given by the Dirac-delta function. We make three standard identifiability assumptions for causal effect estimation: positivity, consistency, and unconfoundedness [e.g., 2, 27]. Finally, we consider an arbitrary machine learning model $\phi$ to predict the potential outcomes. Hence, we define the outcome prediction function as $\phi:\mathcal{X}\times\mathcal{A}\rightarrow\mathbb{R}$, $\phi(X,A)\mapsto Y$. We assume the dose-response curve to be sufficiently Hölder-smooth. This is common in settings with continuous treatments [e.g., 41, 46].
Our objective: In this work, we aim to derive conformal prediction intervals $C(X_{n+1},\Diamond)$ for the prediction of a potential outcome $Y_{n+1}(\Diamond)$ of a new data point under either hard, $\Diamond=a^{*}$, or soft intervention, $\Diamond=A^{*}(X_{n+1})\sim\tilde{\pi}(a\mid X_{n+1})$. The derived intervals are called valid for any new exchangeable sample $X_{n+1}$ with non-exchangeable intervention $\Diamond$, i.e., for $\Diamond\in\{a^{*},A^{*}(X_{n+1})\}$ and significance level $\alpha\in(0,1)$
$P(Y_{n+1}(\Diamond)\in C(X_{n+1},\Diamond))\geq 1-\alpha.$ (1)
Of note, our CP method can be used with an arbitrary ML model $\phi$ to predict the potential outcomes.
In CP, the interval $C$ is constructed based on so-called non-conformity scores [53], which capture the performance of the prediction model $\phi$. For example, a common choice for the non-conformity score is the residual of the fitted model $s(X,A,Y)=\left|Y-\phi(X,A)\right|$, which we will use throughout our work. For ease of notation, we define $S_{i}:=s(X_{i},A_{i},Y_{i})$.
@@
Second, the propensity score $\pi$ is commonly unknown in observational data and, therefore, must be estimated, which introduces additional uncertainty that one must account for when constructing CP intervals ($\rightarrow$ Challenge  b). Crucially, existing coverage guarantees [e.g., 52, 50] do not hold in our setting. Instead, we must derive new intervals with valid coverage under propensity shift.
In the following, we address the above propensity shift by performing a calibration conditional on the shift induced by the intervention, which allows us then to yield valid prediction intervals with significance level $(1-\alpha)$ for potential outcomes of a specific hard or soft intervention. We emphasize that the extension to intervals for causal effects is straightforward, in that one combines the intervals for each potential outcome under a certain treatment and without treatment, so that eventually arrives at CP intervals for the individual treatment effect (ITE). Details are in Supplement A.
@@ 4 CP intervals for potential outcomes of continuous treatments
Recall that intervening on test data breaks the necessary exchangeability assumption, i.e., the guaranteed coverage of at least $(1-\alpha)$. Therefore, we now construct CP intervals where we account for a (potentially unknown) propensity shift in the test data induced by the intervention.
@@
    Example: A doctor prescribes a medication to a new patient. Instead of prescribing the same dosage as he would have prescribed to a similar patient in the past, the doctor is interested in the potential health outcome when increasing (or decreasing) the original dosage by an amount $\Delta_{A}$.
@@
    Unknown propensity score (see Section  4.2): In observational data, the propensity score is unknown. Therefore, we usually assess the effect of hard interventions, i.e., $a^{\ast}$. Here, we face additional uncertainty from propensity score estimation ($\rightarrow$  Challenge  b). [^4] Throughout our main paper, we focus on the setting of hard interventions. In some cases, it might also be of interest to perform soft interventions on the estimated propensity score. We provide derivations for this setting in Supplement A. Example: In our running example, a patient sees a doctor who has never prescribed the respective medication and thus will base the decision on observational data (electronic health records), which was collected under a different, unknown treatment policy (e.g., from another physician). Thus, the prescribed intervention (dosage) cannot be expressed in terms of the policy in the observational data.
@@
Second, our CP method will build upon ideas from so-called split conformal prediction [40, 53], yet with crucial differences. In our methods, the calibration step differs from the standard procedure in that we conditionally calibrate the non-conformity scores depending on the tilting function $f$ to achieve marginal coverage for the interventional – and thus shifted – data.
High-level outline: Our derivation in Sections 4.1 and 4.2 proceed as follows. Following [18, 43], we reformulate split conformal prediction as an augmented quantile regression. Let $S_{i}$ represent the non-conformity score of the sample $(X_{i},A_{i},Y_{i})$ for $i=m+1,\ldots,n$ of the calibration dataset and $S_{n+1}=S$ an imputed value for the unknown score of the new sample. We define
$\hat{\theta}_{S}:=\arg\min_{\theta\in\mathbb{R}}\frac{1}{n-m}\left(\sum_{i=m+1}^{n}l_{\alpha}(\theta,S_{i})+l_{\alpha}(\theta,S)\right),$ (3)
@@
Of note, $\hat{\theta}_{S}$ is an estimator of the $(1-\alpha)$-quantile of the non-conformity scores [32, 47]. Using $\hat{\theta}_{S}$, we construct the CP interval with the desired coverage $(1-\alpha)$. However, the interval is only valid for exchangeable data. Quantile regression might yield non-unique solutions that can depend on the indices of the scores [18], so we later restrict the analysis to solvers invariant to the data ordering. [^5] We note that commonly used solvers, such as interior point solvers, are invariant to the data ordering.